OpenStudy (anonymous):

would 3x-7/12 (12 is below both numbers) be 3x- 7/12 instead 4x-7/12

hero (hero):

?

OpenStudy (anonymous):

its a distributive property the question is 3x-7/12

hero (hero):

\[\frac{3x - 7}{12} = \frac{3x}{12} - \frac{7}{12} = \frac{x}{4} - \frac{7}{12}\]

OpenStudy (anonymous):

Thx